Three 30-ohm resistors are connected in parallel across a 60 – volt source.

   a) What is the effective resistance of the circuit?

 

 

 

 

 

   b) Find the total current drawn from the source.

 

 

 

 

 

   c) Find the current through each resistor.

 



1
Expert's answer
2021-04-30T11:26:30-0400

"\frac{1}{R_T} = \frac {1}{30} + \frac {1}{30} + \frac {1}{30} = \frac{1}{10} = 0.1"

"R_T =\frac{1}{0.1} = 10ohms"


b.)"I =\frac {V}{R}"

Where

V= Voltage

R= Resistance

I= current

"I =\frac {60}{10}= 6Amps"


c.) "I =\frac {V}{R}"

"IR_1 = \frac{60}{30} = 2A"

"IR_2 = \frac{60}{30} =2A"

"IR_3 = \frac{60}{30} = 2A"